Evaluate:




HORIZONTAL MOTION




1. The motion of a particle along the x -axis is described by the equation




m v = (3 s2) t + (0.5 m) t2 s2




where x is in meters, and t is in seconds. Determine the following:




a. Position of the particle 3 seconds after starting




b. Displacement from 2 sec to 5 sec c. Average velocity during first 5 seconds d. Velocity at t = 4secs




UARM:




e. Average acceleration during the second 3 seconds




f. Acceleration at t = 5 secs g. Time the particle moves at 20m/s




h. The acceleration and velocity at start of motion 1. A car is moving at uniform acceleration. Its velocity

1
Expert's answer
2022-02-23T08:31:56-0500

"x(t)= 3t +0.5t^2"


a. Position of the particle 3 seconds after starting

"x(3)= 3*3 +0.5*3^2=13.5\:\rm m"

b. Displacement from 2 sec to 5 sec

"x(5)-x(2)=(3*5+0.5*5^2)\\
-(2*5+0.5*2^2)=15.5\:\rm m"

c. Average velocity during first 5 seconds

"x(5)= 3*5 +0.5*5^2=27.5\:\rm m"

"v_{\rm ave}=\frac{x(5)}{t}=\frac{27.5}{5}=5.5\:\rm m/s"

d. Velocity at t = 4sec

"v=x'=3+t"

"v(4)=3+4=12\:\rm m/s"